OpenStudy (anonymous):

consider the set of all continuous functions with operations of pointwise addition and scalar multiplication, having domain [0, 1]. is this set a vector space? explain.

OpenStudy (anonymous):

check the axioms one at a time that is the point of this exercise

OpenStudy (anonymous):

is it clear what you have to check?

OpenStudy (anonymous):

i am some what confused on what to do since it is a domain and not a vector

OpenStudy (anonymous):

not sure what exactly you mean "the domain is not a vector" the continuous functions are the vectors

OpenStudy (anonymous):

you have to check the axioms for a vector space: \(x+y=y+x\) for example where in this case \(f+g=g+f\) where \(f, g\) are functions continuous on \([0,1]\)

OpenStudy (anonymous):

this is true because by definition \((f+g)(x)=f(x)+g(x)=g(x)+f(x)=(g+f)(x)\)

OpenStudy (anonymous):

you also have to check that \(0\in V\) which it is, it is the zero function i.e. \(0(x)=0\) for all \(x\) is continuous on \([0,1]\)

OpenStudy (anonymous):

most of what you need is inherited from the properties of real numbers, and the fact that addition is point wise. check all the axioms one at a time, you will see it